in 2001 18.2 milion households had a dvd player. If theis population grows at 25% annually, how many households will have a dvd player in 2006 ( 5 years later)? a = p(1+r) to the t power
Answer by checkley71(8403)

A=P(1+R)^T
A=18.2(1+.25)^5
A=18.2(1.25)^5
A=18.2*3.05
A=55.54 MILLION WILL BE THE POPULATION OF DVD PLAYERS IN 5 YEARS.